The length of a certain kind of Colorado brook trout is normally distributed with a mean of 13.1 inches and a standard deviation of 1.5 inches.

What minimum size limit should the Department of Natural Resources set if it wishes to allow people to keep 80 percent of the trout they catch?

Answer by Boreal(15235)

20th percentile is a z of -0.8416
z=(x-mean)/sd
-0.8416=(x-13.1)/1.5
-1.26=x-13.1
x=11.84 inches.